OpenStudy (anonymous):

Factor: 6x^2n + 7x^n - 3 A^2 + 2(A)(B) - B^2?

OpenStudy (anonymous):

If you use substitution you will see the answer more easily. Start by letting u=x^2 then rewrite the equation in terms of u giving \[6u ^{2}+7u -3\] then rewrite as \[6u ^{2}+9u -2u-3=3u \left( 2u+3 \right)-1(2u+3)\] =(2u+3)(3u-1)